HUBBLE SPACE TELESCOPE DAILY REPORT #2317 PERIOD COVERED: 0000Z (UTC) 02/25/99 - 0000Z (UTC) 02/26/99 Daily Status Report as of 057/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Recovery and SMS Science Intercept Recovery from Safemode Inertial Hold was initiated and the SMS was sucessfully started at 056/0500Z. The science observations started at 056/1103Z. 1.2 Completed STIS/CCD/WF/PC-2 7367 (Multi-Wavelength Imaging of Circumstellar Disks in the Orion Nebula) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make multi-wavelength imaging of circumstellar disks (NGC-1976) seen as silhouettes against the bright emission-line background in the Orion Nebula. With these data one should be able to examine the radial structure of the disks in great detail, and over a wide spatial extent. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 7278 (The Connection Between The Obscuring Torus and the Masing Disk in H_20 Me) The WF/PC-2 was used to investigate the connection between the molecular torus and the emission line gas in the H-2 megamaser NGC-1386. The proposal completed nominally with no reported problems. 1.4 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.5 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD/WF/PC-2 7615 (Determination of the Mass Function of a Globular Cluster Through Gravitational Microlensing) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make gravitational microlensing observations of the large globular cluster NGC-6656 located in Sagittarius in order to study its mass function. Gravitational microlensing offers a unique and direct method of determining the masses of stars and other compact objects down to planetary masses. A bias and a dark calibration observation were also taken. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 7465 (White Dwarf Distance and Precision Age for Globular Clusters) The WF/PC-2 was used to make observations of white dwarf star WD1713+659 which is located in globular cluster 47-Tuc. The observed cooling sequences for the clusters is to be fit to the corresponding FCSs (Fiducial Cooling Sequences) to determine accurate cluster distance moduli and absolute ages. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD/MA2 7290 (C-IV In Extensive Gaseous Halos of QSO Lyman-Alpha Absorbing Galaxies) The Space Telescope Imaging Spectrograph (MA2) was used to make observations of the QSO, Lyman-Alpha, absorbing galaxy PKS0405-12. Detection of absorption features would imply that the halo gas of this galaxy has been enriched by some sort of galactic fountain (in place of a halo star formation) and would also determine up to which radius this process has operated. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset once. The SSA transmitter was turned on and off per ROP IC-2. Gyro-3, which is currently operating in the shadow mode, had a saturation error at 057/041630Z for one 12-second sample during a T2 slew and near a night/day transition. HSTAR 7018 was written. The Gyro-3 drift rate bias trend data shows another decrease. The rate of change over the last two days is 30.89 arcseconds/hour/day (was 34.49). the trend still exceeds the desired 10.8 arcseconds/hour/day. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH